OpenStudy (shaeelynn):

rationalize denominators to simplify radical expressions 4th root 2/4th root 5

OpenStudy (shaeelynn):

\[\frac{ \sqrt[4]{2} }{ \sqrt[4]{5} }\]

OpenStudy (agent0smith):

These suck. It's probably easier to turn things into fractional exponents first \[\large \frac{ 2^{1/4} }{ 5^{1/4} }\] You have to multiply the top and bottom by something that'll make the exponent on the 5 into a 1. \[\Large \frac{ 2^{1/4} }{ 5^{1/4} }*\frac{ 5^{3/4} }{ 5^{3/4}}\] why 3/4? because 1/4 + 3/4 is...

OpenStudy (shaeelynn):

1

OpenStudy (agent0smith):

\[\Large \frac{ 2^{1/4} }{ 5^{1/4} }*\frac{ 5^{3/4} }{ 5^{3/4}} = \frac{ 2^{1/4} *5^{3/4}}{ 5 }\]

OpenStudy (shaeelynn):

is it \[\frac{ \sqrt[4]{250} }{ 5 }\]

OpenStudy (agent0smith):

Now turn the 1/4 exponents back into 4th root symbols\[\Large \frac{ \sqrt[4]{2} * \sqrt[4]{5^{3}}}{ 5 }=\]

OpenStudy (shaeelynn):

okay so the answer is \[\frac{ \sqrt[4]{250} }{5 }\]

OpenStudy (shaeelynn):

can you help me with another problem?

OpenStudy (agent0smith):

Okay

OpenStudy (shaeelynn):

\[(\sqrt{y}+\sqrt{2})(\sqrt{y}-7\sqrt{2})\]

OpenStudy (agent0smith):

Expand it like you would any expression like that.

OpenStudy (shaeelynn):

what?

OpenStudy (agent0smith):

Like any expression of that form eg. (x+2)(x-6)

OpenStudy (shaeelynn):

ohh distribute

OpenStudy (shaeelynn):

i forgot how to distribute radicals

OpenStudy (agent0smith):

Remember drawing the lines, from each term to each other term in the parentheses... so you don't forget anything

OpenStudy (shaeelynn):

yes i remember that part

OpenStudy (shaeelynn):